# Submitted toTopology Proceedings

### v2

@@ Abstract.
In this follow-up paper to [10] we investigate the global topology and geometry of dynamical systems $\dot{x}=F(x)$ with entire vector field $F$ by exploiting and extending the local structure of finite elliptic decompositions. We prove topological properties of the basins of centers, nodes, and foci, while excluding isolated equilibria at the boundaries of the latter two. We propose a definition of global elliptic sectors and introduce the notion of a sector-forming orbit allowing global topological analysis. The structure of heteroclinic regions between two equilibria is characterized.
@@
Holomorphic dynamical system, global phase portrait, basin of attraction/stability, global elliptic sector, heteroclinic region
@@
Acknowledgement: We thank Jörn Dietrich for mathematical discussions and collaborative solution-finding.
@@ 1. Introduction
In our previous paper [10], we focused on the local topological properties of both simple and higher-order equilibria to set the stage for investigating the global phase portrait of holomorphic vector fields in this follow-up paper. In order to reveal the relation between local and global dynamics, we utilize the geometric structure of finite elliptic decompositions to explore the "stability regions" of equilibria, global elliptic sectors, and heteroclinic regions. This is the next step towards studying holomorphic and meromorphic flows of vector fields derived from complex functions pertinent to number theory, such as the Riemann $\zeta$- and $\xi$-function, cf. [5, 6, 11, 21]. The boundary-forming orbits (separatrices) of these "stability regions" are particularly interesting since they divide the complex plane into regions that exhibit completely distinct topological behavior, reminiscent of phase transitions. In previous works, we have been analyzing normally attracting/repelling (slow) invariant manifolds of equilibria in the context of multi scale dynamical systems, cf. [8, 12, 13, 14, 15, 16, 17], which, in the 2-dimensional planar case, are usually separatrices dividing the phase space into topological basins.
In the following, we provide an outline of our paper. We consider the special case of entire vector fields in real time parameterization, i.e. a holomorphic dynamical system of the form
@@
with $F\in\mathcal{O}(\mathbb{C}{})$. [^1] In the context of dynamical systems, we use the notation of orbits and flows given in [10, Chapter 2]. A definition of limit sets, cf. [2, §4] and [20, Chapter 8.4], as well as some basic results can also be found in [10, Chapter 2]. We start with the analysis of the basin of a center, cf. [4, Theorem 4.1] and [3, Theorem 3.3], and provide detailed proofs of several topological properties. In particular, we detail the proof in [3, Theorem 3.3] for the openness of the basin of a center.
Subsequently, we examine the topological structure of equilibria of order $m\geq 2$. Since a precise definition of global elliptic sectors is missing in the literature222cf. [2, Chapter VIII, §17], [4, Chapter 4], [7, Chapter 1.5] and [19, Chapter 2.11]., we propose an appropriate characterization by introducing the concept of sector-forming orbits. We demonstrate that the intuitive idea of the geometric structure of a global elliptic sector is indeed rigorously valid. We support the proof of the openness with geometric visualization, cf. Figure 1, and point out the conditions under which the local structure of a finite elliptic decomposition can be transferred to the global phase portrait, an issue already addressed in [10, Example 4.7, Example 4.8].
@@
Building on these results, we analyze the topology of the basin of nodes and foci. Notably, the authors of [4, Theorem 4.3] explicitly do not exclude the scenario of an isolated equilibrium on the boundary of the basin – specifically, an equilibrium that is not attached to an orbit lying outside the basin. We address this aspect in detail in Proposition 4.5.
@@
Finally, we define and investigate heteroclinic regions between two equilibria, observing properties similar to the preceding cases. This leads to the exclusion of the existence of a cycle in the phase space formed by equilibria and heteroclinic orbits connecting these equilibria, cf. Corollary 5.7.
@@ 2. Basin of centers
In this chapter, we treat the case of a center. We provide more detailed complete proofs of topological results given in [4, Theorem 4.1] and [3, Theorem 3.3]. In all considerations of this chapter we consider an entire function $F\in\mathcal{O}{}(\mathbb{C}{})$, $F\not\equiv 0$, and a center $a\in\mathbb{C}{}$ of (1.1).
@@
We define the basin $\mathcal{V}$ of $F$ in $a$ as
@@
[10, Definition 3.1] ensures $a\in\mathring{\mathcal{V}}$, i.e. $\mathring{\mathcal{V}}\not=\emptyset$. Let $y\in\mathcal{V}$ and $z\in\Gamma(y)$. Then $\Gamma(y)=\Gamma(z)\subset\mathcal{V}$. Thus $\mathcal{V}$ is flow-invariant. Every $y\in\operatorname{Int}(\Gamma(x))$ either satisfies $y=a$, or, by [10, Theorem 5.1], is a periodic orbit with $a$ in its interior. This proves equation (2.1). By [22, Lemma 6.4], it follows that also $\partial\mathcal{V}$ is flow-invariant. ∎
@@
It holds $\partial\mathcal{V}\cap F^{-1}(\{0\})=\emptyset$, i.e. there are no equilibria on the boundary of $\mathcal{V}$.
@@
The case $\partial\mathcal{V}=\emptyset$ is trivial, so let $x\in\partial\mathcal{V}$ and assume this is an equilibrium. For all $\varepsilon>0$ there exists $y\in\mathcal{B}_{\varepsilon}(x)\cap\mathcal{V}$ such that $\Gamma(y)$ neither has only $x$ in its limit sets nor is a periodic orbit with $x$ in its interior. Note that $x\not=a$, since $a\in\mathring{\mathcal{V}}$. Hence, $x$ cannot be a center, focus or node. By [10, Theorem 4.4], $x$ has a FED. In particular, by [10, Definition 4.1], there exists $r>0$ such that $x\in w_{+}(\Gamma(z))\cup w_{-}(\Gamma(z))$ for all $z\in\mathcal{B}_{r}(x)$. But no orbits in $\mathcal{B}_{r}(x)\cap\mathcal{V}\not=\emptyset$ tend to $x$, i.e. this is a contradiction to [10, Definition 4.1 c) (ii)]. Thus, $x$ cannot be an equilibrium. ∎
@@
$\mathcal{V}$ is open.
@@
Suppose there exists $x\in\mathcal{V}\cap\partial\mathcal{V}\not=\emptyset$, $x\not=a$. Then $\Gamma(x)\subset\mathcal{V}\cap\partial\mathcal{V}$, cf. Proposition 2.2. In particular, $\Gamma(x)$ is a periodic orbit with $a\in\operatorname{Int}(\Gamma(x))\not=\emptyset$. Suppose there also exists $y\in\partial\mathcal{V}\setminus\Gamma(x)\not=\emptyset$. The set $\mathbb{C}\setminus\Gamma(x)=\operatorname{Int}(\Gamma(x))\cup\operatorname{Ext}(\Gamma(x))$ is open. If $y\in\operatorname{Int}(\Gamma(x))$, there exists $r>0$ such that $\mathcal{B}_{r}(y)\subset\operatorname{Int}(\Gamma(x))$. By equation (2.1), we get $\mathcal{B}_{r}(y)\subset\mathcal{V}$, i.e. $y\not\in\partial\mathcal{V}$. This a contradiction. Furthermore, if $y\in\operatorname{Ext}(\Gamma(x))$, there exists again an $r>0$ such that $\mathcal{B}_{r}(y)\subset\operatorname{Ext}(\Gamma(x))$. Since $y$ lies on the boundary of $a$, there exists $z\in\mathcal{B}_{r}(y)\cap\mathcal{V}$, i.e. $\Gamma(z)$ is a periodic orbit with $a\in\operatorname{Int}(\Gamma(z))$ and $\Gamma(z)\not=\Gamma(x)$. Since $\Gamma(z)$ is path-connected and $z\in\operatorname{Ext}(\Gamma(x))$, we must have $x\in\operatorname{Int}(\Gamma(z))$. By equation (2.1) and openness of $\operatorname{Int}(\Gamma(z))$, we derive the contradiction $x\in\mathring{\mathcal{V}}$. Thus, $y\not\in\operatorname{Int}(\Gamma(x))\cup\operatorname{Ext}(\Gamma(x))$, which is also a contradiction. We get $\partial\mathcal{V}\setminus\Gamma(x)=\emptyset$, i.e. the boundary of $a$ consists exactly of the unique outermost periodic orbit $\Gamma(x)$ and $\mathcal{V}=\overline{\operatorname{Int}(\Gamma(x))}$ must be compact.
By equation (2.1), $\Gamma(x)$ is not isolated. Hence we can apply a result on limit cycles of analytic dynamical systems, [2, p. 203, §12.1, Lemma 1], to deduce the existence of $\rho>0$ such that all orbits through the set $U:=\{y\in\mathbb{C}{}:\text{dist}(\Gamma(x),y)<\rho\}\cap\operatorname{Ext}(\Gamma(x))$ are periodic. Reduce $\rho$ such that $\overline{\operatorname{Int}(\Gamma(x))}\cup U$ is simply connected and $U\cap F^{-1}(\{0\})=\emptyset$. By continuity of the flow, cf. [19, Chapter 2.4, Theorem 4], we deduce the existence of at least one periodic orbit $\Gamma\subset U$. By choice of $\rho$, we get $\operatorname{Int}(\Gamma)\subset U\cup\overline{\operatorname{Int}(\Gamma(x))}$ and thus $a\in\operatorname{Int}(\Gamma)$, cf. [2, §11, Corollary 1] and [3, Theorem 2.4]. This is a contradiction to $\Gamma\subset\mathbb{C}{}\setminus\mathcal{V}$. Hence the assumed $x$ cannot exist and $\mathcal{V}$ is open. ∎
@@
The interior of any periodic orbit is clearly connected. Hence, by Proposition 2.2, $\mathcal{V}$ can be interpreted as union of connected sets, all having the point $a$ in common. Thus, also the union must be connected, cf. [18, Theorem 23.3]. By [23, Proposition 12.25], $\mathcal{V}$ is even path-connected.
Let $L\subset\mathcal{V}$ be a loop (closed path). By openness of $\mathcal{V}$, cf. Theorem 2.4, for all $x\in L$ there exists a $y_{x}\in\operatorname{Ext}(\Gamma(x))\cap\mathcal{V}\not=\emptyset$ such that $x\in\operatorname{Int}(\Gamma(y_{x}))\subset\mathcal{V}$. Hence, the system $\left\{\operatorname{Int}(\Gamma(y_{x})):x\in L\right\}$ is an open cover of $L$. Since $L$ is compact, there exist finitely many points $x_{1},\ldots,x_{\ell}\in L$, $\ell\in\mathbb{N}{}$, such that
$L\subset\bigcup_{k=1}^{\ell}\operatorname{Int}(\Gamma(y_{x_{k}}))\subset\mathcal{V}\text{.}$
Since $\ell$ is finite, $y\in\left\{y_{x_{k}}:1\leq k\leq\ell\right\}$ must exist such that $\Gamma(y)$ is the outermost periodic orbit having the other $\ell-1$ periodic orbits in its interior. By equation (2.1), we conclude $L\subset\operatorname{Int}(\Gamma(y))\subset\mathcal{V}$ with $y\in\mathcal{V}$. Since $\operatorname{Int}(\Gamma(y))$ is clearly simply connected, $L$ can be continuously (homotopically) deformed to a point in $L\subset\mathcal{V}$. Since $\mathcal{V}$ is already path-connected, it follows that $\mathcal{V}$ is simply connected. ∎
Theorem 2.6.
All orbits on the boundary of $a$ are unbounded. $\mathcal{V}$ is unbounded.
@@
If the boundary of $a$ is empty, then $\mathcal{V}$ is open, closed and nonempty, i.e. $\mathcal{V}=\mathbb{C}{}$ and thus unbounded. Assume $\partial\mathcal{V}\not=\emptyset$ and let $x\in\partial\mathcal{V}$. Suppose that $\Gamma(x)\subset\partial\mathcal{V}$ is bounded. By Proposition 2.3, we get $w_{\pm}(\Gamma(x))\cap F^{-1}(\{0\})=\emptyset$. Hence, by [10, Corollary 5.2], $\Gamma(x)$ must be periodic. By [10, Theorem 5.1], $\Gamma(x)$ has exactly one equilibrium, a center $\tilde{a}$, in $\operatorname{Int}(\Gamma(x))$. If $\tilde{a}\not=a$, $x\in\partial\mathcal{V}$ would be a point in the basin in $\tilde{a}$, which is a contradiction to the openness of this basin, cf. Theorem 2.4. If $a=\tilde{a}$, we get $x\in\mathcal{V}$, which is again a contradiction, since $\mathcal{V}\cap\partial\mathcal{V}=\emptyset$. Thus, $\Gamma(x)$ and $\mathcal{V}$ must be unbounded. ∎
@@ 3. Global elliptic sectors
In all claims and arguments of this chapter, we consider an entire function $F\in\mathcal{O}{}(\mathbb{C})$, $F\not\equiv 0$, and an equilibrium $a\in\mathbb{C}$ of (1.1) of order $m\in\mathbb{N}\setminus\{1\}$. By applying [10, Proposition 4.3] and [10, Theorem 4.4], (1.1) has a finite elliptic decomposition (FED) of order $22m-$ 2 in $a$, where all local elliptic sectors have adjacent definite directions given by $\mathcal{E}(F,m)$. Moreover, we use the function $\lambda:\mathcal{E}(F,m)\to\{-1,1\}$, $\lambda(\theta):=\cos(\operatorname{arg}(F^{(m)}(a))+\theta(m-1))$ to determine the time direction (positive or negative) in which the orbits tend to $a$ in the FED, cf. [10, Proposition 4.3 (iii)] and [10, calculation of $\lambda_{i}$ on p. 11].
In the following, we will extend the analysis of the local structure described in [10, Definition 4.1 c)] to a global point of view and provide detailed complete proofs of the topological results given in [4, Theorem 4.2]. To do this, we need at first a precise definition of a global elliptic sector.

### v3

@@ Abstract.
In this follow-up paper, we investigate the global geometry and topology of dynamical systems $\dot{x}=F(x)$ with entire vector field $F$, building on and constructively extending the local structure of simple and higher-order equilibria. We provide a step-by-step analysis to reveal topological properties of the basins of centers, nodes, and foci, while excluding isolated equilibria at the boundaries of the latter two. We propose a definition of global elliptic sectors and introduce the concept of sector-forming orbits based on the geometry within a finite elliptic decomposition of multiple equilibria. Finally, we characterize the structure of heteroclinic regions connecting two equilibria.
@@
Holomorphic dynamical system, entire vector field, center, period annulus, focus, node, basin of attraction/stability, elliptic sector, canonical region
@@
Acknowledgement: We thank Jörn Dietrich for mathematical discussions and collaborative solution-finding. We thank the referee for helpful comments, suggestions, and additional literature references.
@@ 1. Introduction
In our previous work [19], we examined the local structure of simple and higher-order equilibria in holomorphic vector fields, i.e. in holomorphic dynamical systems of the form
@@
with $F\in\mathcal{O}(\mathbb{C}{})$. Building on these local insights, this follow-up paper shifts the perspective to the global phase portrait of entire vector fields and develops an understanding of how local structures influence and shape the global geometry of (1.1). Throughout this paper, we use the terminology and notion of orbits and flows in the context of dynamical systems and the Jordan curve Theorem from [19, Chapter 2]. In particular, if $\Gamma\subset\mathbb{C}{}$ is a closed Jordan curve, we denote the two connecting components resulting from the Jordan curve Theorem by $\operatorname{Int}(\Gamma)$ (the bounded interior of $\Gamma$) and $\operatorname{Ext}(\Gamma)$ (the unbounded exterior of $\Gamma$).
As examined in [19], each equilibrium of (1.1) falls into one of the following categories:
  • (i)
    A center (simple equilibrium), where all orbits in a neighborhood are closed periodic orbits surrounding the center.
  • (ii)
    A focus or node (simple equilibrium), which is attracting or repelling such that all nearby orbits tend to the equilibrium either in positive (attracting) or negative (repelling) time.
  • (iii)
    A multiple equilibrium of order $m\in\mathbb{N}\setminus\{1\}$, possessing a finite elliptic decomposition of order 2m22m-2. This geometric structure is defined in [19, Definition 4.1] and illustrated in Section 3.1.
With this in mind, we focus on the following two central questions to reveal the relation between local and global dynamics:
  • (A)
    How can we globally define the "region of influence" of an equilibrium based on its known local structure, and what are the topological properties of the resulting set?
  • (B)
    How can these "regions of influence" and their geometric structure be embedded into the existing theory of canonical (strip, annular, radial, and spiral) regions in general smooth planar dynamical systems, cf. [8, 9, 10, 11]?
For each of the three cases (i)–(iii) described above, we provide a step-by-step analysis to address questions (A) and (B). In addition, we emphasize the significance and specific influence of the holomorphy of the vector field in (1.1) on the possible geometric configurations of the global phase space. This becomes apparent both in the results stated in theorems and in the arguments within proofs. In particular, the central tool is [19, Corollary 5.1], which can be formulated in the case of entire vector fields as follows:
Theorem 1.1 (Poincaré-Bendixson for Holomorphic Flows).
Let $F\in\mathcal{O}{}(\mathbb{C})$, $F\not\equiv 0$, be an entire function and $K\subset\mathbb{C}$ compact with $\xi\in K$ such that $\Gamma_{+(-)}(\xi)\subset K$. Then either $\Gamma(\xi)$ is a periodic orbit with exactly one equilibrium, a center, in its interior, or the positive (negative) limit set of $\Gamma(\xi)$ consists of exactly one equilibrium. Additionally, if $\Gamma(\xi)$ is periodic, the interior of $\Gamma(\xi)$ (except for the center) is filled entirely with periodic orbits all having the center in its interior.
The analysis of questions (A) and (B) also sets the stage for studying holomorphic and meromorphic flows of vector fields derived from complex functions pertinent to number theory, such as the Riemann $\zeta$- and $\xi$-function [6, 7, 20, 21, 31] and their polynomial approximations. Some ideas are motivated by previous works, in which we analyzed normally attracting or repelling (slow) invariant manifolds of equilibria within the context of multiscale dynamical systems [17, 22, 23, 24, 25, 26, 27]. In the following, we provide an outline of our paper.
We start with a short chapter about the center basin, also called period annulus. We use the well-known theory about closed orbits examined in [2, 9, 11] in order to provide complete proofs of several topological properties that are already stated in [5, Theorem 4.1] and [4, Theorem 3.3].
Subsequently, as the main focus of this paper, we examine the topological structure of equilibria of order $m\geq 2$. Since a precise definition of a global elliptic sector is lacking in the literature, cf. [2, 5, 13, 14, 16, 29], we introduce the concept of sector-forming orbits to provide an appropriate global characterization. We show that the intuitive notion of the geometric structure of a global elliptic sector can indeed be made rigorous. We conclude this section by showing that, in the sense of Markus and Neumann [8, 9], it naturally forms a strip canonical region, thus addressing question (B).
@@
Building on these results, we analyze the topology of the basin of nodes and foci. The authors of [5, Theorem 4.3] explicitly do not exclude the scenario of an isolated equilibrium on the boundary of the basin. We address this aspect in detail in Proposition 4.4 with a rigorous proof in the appendix.
@@
Finally, we define and investigate heteroclinic regions as one geometric possibility between two adjacent elliptic sectors, observing topological properties similar to the preceding cases. This leads to the exclusion of the existence of a cycle in the phase space formed by equilibria and heteroclinic orbits connecting these equilibria, cf. Corollary 5.6. We conclude this paper with an example that illustrates the occurrence of several elliptic sectors and heteroclinic regions together with their separatrices in the phase space.
Throughout the paper, we selectively focus on explicit constructions that extend local structures to global ones, without a priori resorting to potentially known or conjectured global results, because we consider constructive proofs to offer a fruitful setting for developing general approaches to questions of type (A) and (B). Nevertheless, we aim to contextualize our framework within previous research from different but related perspectives.
@@ 2. Basin of centers
Throughout this chapter, we consider an entire function $F\in\mathcal{O}{}(\mathbb{C}{})$, $F\not\equiv 0$, and a center $a\in\mathbb{C}{}$ of (1.1).
@@
We define the center basin or period annulus $\mathcal{V}$ of $F$ in $a$ as
@@
A well-established theory exists regarding the local and global geometry of the period annulus $\mathcal{V}$. Markus and Neumann, cf. [8, 9, 11], provide a thorough analysis of annular (parallel) canonical regions. In the following, we highlight which of the known properties extend to the holomorphic setting and examine how holomorphy further refines both the global geometry and the proof techniques. This approach leads to complete proofs of the topological results stated in [5, Theorem 4.1] and [4, Theorem 3.3].
@@
Definition 3.1 in [19] ensures $a\in\mathring{\mathcal{V}}\not$$\emptyset$. The flow-invariance of $\mathcal{V}$ and $\partial\mathcal{V}=\overline{\mathcal{V}}\cap\overline{\mathbb{C}\setminus\mathcal{V}}$ follow from [32, Lemma 6.4]. By using the holomorphy of $F$, we can apply Theorem 1.1 to conclude that all orbits in $\operatorname{Int}(\Gamma(x))$, with $x\in\mathcal{V}$, are periodic with $a$ in the interior, which shows (2.1). ∎
This Proposition shows that $\mathcal{V}$ is indeed a parallel annular region, cf. [11, p. 90]. In general, equation (2.1) does not hold for arbitrary planar dynamical systems. In fact, there are examples where one or more limit cycles appear within closed orbits, with spiraling behavior in a neighborhood of the limit cycle. However, holomorphy rules out the existence of limit cycles and similar phenomena within periodic orbits, cf. [4, Theorem 3.2].
The following result demonstrates that, once again due to holomorphy of $F$, no equilibria can lie on the boundary of $a$.
@@
It holds $\partial\mathcal{V}\cap F^{-1}(\{0\})=\emptyset$.
@@
The case $\partial\mathcal{V}=\emptyset$ is trivial, so let $x\in\partial\mathcal{V}$ and assume this is an equilibrium. For all $\varepsilon>0$ there exists $y\in\mathcal{B}_{\varepsilon}(x)\cap\mathcal{V}$ such that $\Gamma(y)$ neither has only $x$ in its limit sets nor is a periodic orbit with $x$ in its interior. Note that $x\not=a$, since $a\in\mathring{\mathcal{V}}$. Hence, $x$ cannot be a center, focus or node. By using the holomorphy of $F$, we can apply [19, Theorem 4.4] to conclude that $x$ has a FED. In particular, by [19, Definition 4.1], there exists $r>0$ such that $x\in w_{+}(\Gamma(z))\cup w_{-}(\Gamma(z))$ for all $z\in\mathcal{B}_{r}(x)$. But the orbits in $\mathcal{B}_{r}(x)\cap\mathcal{V}\not$$\emptyset$ do not tend to $x$, contradicting the property c) (ii) in [19, Definition 4.1]. Thus, $x$ cannot be an equilibrium. ∎
@@
$\mathcal{V}$ is open and unbounded. All orbits on $\partial\mathcal{V}$ are unbounded.
@@
Since $\mathcal{V}$ consists of closed periodic orbits, we can use the already known theory about annular regions of analytic planar dynamical systems. In [2, §12.1, p. 203], it is shown that $\partial\mathcal{V}$ consists of equilibria as well as separatrices of saddle points. However, the holomorphy of $F$ prevents the existence of saddles, cf. [19, Theorem 3.2]. Hence, by Proposition 2.3, $\partial\mathcal{V}$ consists only of unbounded separatrices tending to infinity in both time directions. In particular, each of these separatrices connects two saddles at Infinity, cf. [12], and is in particular not a periodic orbit, i.e. $\mathcal{V}\cap\partial\mathcal{V}=\emptyset$. ∎
We have thus verified that $\mathcal{V}$ is a canonical region whose boundary consists of separatrices, cf. [9, Theorem 5.2]. Furthermore, there exists a global smooth transversal in $\mathcal{V}\setminus\{a\}$, that is, an open smooth arc crossing each closed orbit in $\mathcal{V}$ exactly once. The definition of annular canonical regions leads immediately to the following result. However, we present a alternative short proof by applying a version of the Seifert–Van Kampen theorem, combined with the results obtained above.
@@
The interior of any periodic orbit is clearly connected. Hence, by Proposition 2.2, $\mathcal{V}$ can be interpreted as union of connected sets, all having the point $a$ in common. Thus, also the union must be connected, cf. [28, Theorem 23.3]. By [33, Proposition 12.25], $\mathcal{V}$ is even path-connected. By openness of $\mathcal{V}$ we get
$\mathcal{V}=\bigcup_{\mathclap{x\in\mathcal{V}}}\operatorname{Int}(\Gamma(x))\text{.}$
All the sets $\operatorname{Int}(\Gamma(x))$, with $x\in\mathcal{V}$, are open and simply connected. Hence, by applying the Seifert-van Kampen theorem for infinite open covers, cf. [15, Theorem 1.20], we conclude that also $\mathcal{V}$ is simply connected. ∎
We conclude this chapter with a lemma that will play an important role at several points in what follows.
Lemma 2.6.
Let $A\subset\mathbb{C}$ be a flow-invariant set that admits no closed orbits. Then $\mathcal{V}\cap\partial A=\emptyset$.
@@
If the set $\mathcal{V}\cap\partial A$ was not empty, it would contradict the openness of $\mathcal{V}$, cf. Theorem 2.4. ∎
@@ 3. Global elliptic sectors
In this chapter, we consider an entire function $F\in\mathcal{O}{}(\mathbb{C})$, $F\not\equiv 0$, and an equilibrium $a\in\mathbb{C}$ of (1.1) of order $m\in\mathbb{N}\setminus\{1\}$. Our aim is to define and understand the global structure of the "areas of influence" associated with such an equilibrium. We begin by providing a brief overview of earlier research on this topic.
3.1. Survey of related research
In [2, §17] and [16, Chapter 1.5], the authors describe and analyze several types of curvilinear sectors in general planar dynamical systems, providing detailed proofs of their geometric properties. A global analysis of planar dynamical systems is given in [8, 9, 10, 11], where canonical regions and global transversals are classified. The structure of multiple equilibria is also discussed there. However, these works do not specifically address the special case of holomorphic vector fields in (1.1).
In [13], the authors classify holomorphic flows up to conformal equivalence in a neighborhood of an equilibrium, providing representatives for each equivalence class. This yields several model flows together with their corresponding local phase portraits. A related investigation is carried out in [3], where the authors describe various normal forms depending on the order and residue of the equilibrium. For example, the local structure of an equilibrium $a$ of order $m\in\mathbb{N}\setminus\{1\}$ is conformally equivalent to (1.1) with
$F(x):=x^{n}-cx^{2n-1}\text{,}$
where $c\in\mathbb{C}{}$ is the residuum of $\frac{1}{F}$ at $a$. However, all these analyses concern the local phase portrait.
We have likewise analyzed the local structure of multiple equilibria in our recent paper [19], where we established the existence of a finite elliptic decomposition (FED) of order 2m22m-2 in the multiple equilibrium $a$, cf. [19, Proposition 4.3] and [19, Theorem 4.4]. In particular, we showed that each local elliptic sector of this decomposition has adjacent definite directions given by
$\mathcal{E}(F,m)=\left\{\frac{\ell\pi-\arg(F^{(m)}(a))}{m-1}\mod 2\pi:\ell\in\mathbb{Z}\right\}\subset[0,2\pi)\text{.}$
The geometry of a local elliptic sector is defined in [19, Definition 4.1 c)] and illustrated in Figure 1. A finite elliptic decomposition can be obtained by cyclically copying the geometry in Figure 1 around the equilibrium. In the proof of [19, Theorem 4.4], we established step by step the various geometric objects required for a finite elliptic decomposition without invoking the theory of topological/conformal equivalence.
[…]
Figure 1. Geometrical objects of a local elliptic sector (blue) in a multiple equilibrium $a$. $\Gamma_{1}$ and $\Gamma_{2}$ (black) are the characteristic orbits. $\Lambda_{1}$ and $\Lambda_{2}$ (red) are the transversals or paths without contact. $\Xi=\Gamma(E_{1})$ (green) is the homoclinic orbit. The arrows indicate the direction of the vector field.
With this in mind, we now extend the analysis of the local structure to a global perspective by introducing the notion of a global elliptic sector. We provide detailed complete proofs of the topological results given in [5, Theorem 4.2]. Furthermore, we embed this global elliptic sector as a canonical region into the global theory of Markus and Neumann, cf. [9, 10, 11].
3.2. The geometry of global elliptic sectors
In this chapter, we provide a step-by-step analysis of the global elliptic sector to obtain a complete topological and geometrical characterization. We introduce the notion of the global elliptic sector in an intuitive manner by "blowing up" the local structure in an appropriate way.